Solve for each rational equation. Check your solutions.
√(5x-4=6) The square root of 5x-4=6

Answer by checkley79(3341)

sqrt(5x-4)=6 square both sides.
5x-4=35
5x=36+4
5x=40
x=40/5
x=8 ans.
Proof:
sqrt(5*80-4)=6
sqrt(40-4)=6
sqrt(36)=6
6=6